# Spatiotemporal Density Correction of Multivariate Global Climate Model Projections using Deep Learning

### v2

@@ 3.3 Comparison of SPCDE with CCA and QM
SPCDE CCA QM SPCDE CCA QM
TMAX Wasserstein distance 0.3001 0.9270 0.3073 1.0377 1.3454 1.6735
0.95 quantile 1.2225 0.9202 1.1136 1.4198 1.7515 2.2220
Lag-1 autocorrelation 0.0878 0.6873 0.6340 0.0617 0.6768 0.6261
Spatial correlation 0.0567 0.0737 0.0670 0.0231 0.0653 0.0639
PRCP Wasserstein distance 0.1049 0.5427 0.1256 0.3057 0.8573 0.4780
0.95 quantile 1.7358 2.3285 1.6520 4.9077 4.9544 5.8247
Proportion of zeros 0.1222 0.2111 0.2227 0.0818 0.1056 0.1978
Lag-1 autocorrelation 0.1441 0.4102 0.3539 0.0970 0.2978 0.2362
Spatial correlation 0.1205 0.1281 0.1283 0.0528 0.1401 0.1382
Cross correlation 0.0017 0.3209 0.3209 0.0006 0.1039 0.1039
@@
To evaluate the performance of the SPCDE density correction approach compared to CCA and QM, we provide metrics for predictive accuracy based on the three methods in Table 1. Marginal fits for TMAX and the transformed PRCP are summarized using the Wasserstein distance, while for the remainder of the metrics provided in this section, we evaluate the RMSE. In each case, smaller values indicate more accurate predictions. We note that the SPCDE predictions are better than CCA and QM predictions across the board. The only two instances where SPCDE is not outright better are the 0.95 quantiles for SW TMAX and PRCP. Neither CCA nor QM is clearly better even in this case, with the former having lower RMSE for TMAX, and the latter having a lower RMSE for PRCP. On the other hand, SPCDE is still the best among the three methods at estimating the proportion of zeros in PRCP for both regions. The autocorrelations are also significantly better for SPCDE, since neither of the competing method explicitly accounts for autocorrelation. While SPCDE has smaller RMSE for spatial correlations in every case, the differences are more noticeable for the SE, which has both lower bias and variability in its spatial correlation estimates (see Figure 7). Finally, RMSE in the cross correlations are around 3 three orders of magnitude smaller for SPCDE than competing methods.
@@
Overall, the SPCDE calibrated distributions of TMAX and PRCP have the same characteristics as their observed data counterparts, for statistics of interest. In particular, it is able to very accurately capture the relationship between the two variables, which is crucial as they jointly affect several quantities of practical interest, including droughts, fire risk, and streamflow.
@@ 4 Discussion
In this paper, we develop a new SPCDE approach for the simultaneous density correction of spatiotemporal PRCP and TMAX. A key goal of our study is to maintain the cross correlation between the two variables during density correction. SPCDE leverages the Vecchia approximation to maintain spatiotemporal and inter-variable dependencies, and density estimation is carried out using the deep learning based SPQR method. Our approach is trained across two spatial fields in the Southeast and Southwest US using historical data from 1951–2000, and tested using data from 2001–2014 for the same region. SPCDE is able to outperform QM and CCA, two competing methods, and the joint distribution characteristics of the observational data is preserved in the bias-corrected data.
@@
SPCDE is highly scalable, since the Vecchia approximation allows for fitting models at different locations and months in parallel. Additionally, it ensures that the computational cost is linear in the number of grid cells. In our current implementation, however, we fit separate models for each month so that seasonality does not affect our results. Future work will focus on consolidating the monthly models into a single annual model to reduce computational cost and improve estimation (since significantly more data will be available to train the annual model). We anticipate this can be achieved by including more time-based covariates to account for trends and seasonality, in a way that improves the estimation of autocorrelation in the data. Additionally, we plan to extend SPQR to support a bivariate response variable, which could then be used to directly model the joint distribution of TMAX and PRCP without needing to decompose it into a product of a conditional and a marginal distribution. The ultimate goal of these anticipated extensions is to provide future projections of TMAX and PRCP across the entire contiguous US.
Acknowledgements
@@
This work was supported by a grant from the National Science Foundation (DMS 2152887, CBET 2151651). Hector was also supported by an Internationalization Seed Grant award from the Office of Global Engagement at North Carolina State University.
Appendix A Additional results
In this section, we provide results analogous to Section 3, but focused on in-sample performance. Instead of training the SPCDE on 1951–2000 data, we train it on the entire dataset from 1951–2014. While the focus of this study was the evaluation of SPCDE during the historical period, the eventual goal is future climate projections. In such cases, we would train SPCDE on the entire historical record. The plots and the table provided in this section are identical in all aspects as those provided in Section 3, and only differ in the temporal extent of the data.
Figure 8: Density curves of TMAX and PRCP data in the Southwest (SW) and Southeast (SE) for 1951–2014 data. PRCP results are presented in the log scale.
Figure 9: Monthly 0.95 quantiles of TMAX and PRCP data in the Southwest (SW) and Southeast (SE), aggregated over 1951–2014. Each point corresponds to a particular month and location.
Figure 10: Monthly Proportion of zeros in PRCP data for the Southwest (SW) and Southeast (SE), aggregated over 1951–2014. Each point corresponds to a particular month and location.
Figure 11: Monthly lag-1 autocorrelations of TMAX and PRCP data in the Southwest (SW) and Southeast (SE), aggregated over 1951–2014. Each point corresponds to a particular month and location.
Figure 12: Monthly cross-correlations between TMAX and PRCP in the Southwest (SW) and Southeast (SE), aggregated over 1951–2014. Each point corresponds to a particular month and location.
Figure 13: Annual spatial correlations of TMAX and PRCP data in the Southwest (SW) and Southeast (SE), aggregated over 1951–2014. Each point corresponds to a pair of locations.
@@ Appendix A Additional results
Table 2: Out-of-sample prediction accuracy in the density correction of TMAX and PRCP based on a spatiotemporal SPQR model, compared with canonical correlation analysis (CCA) and quantile mapping (QM). Other than the Wasserstein distance metric, the remaining values correspond to the RMSE between observed and calibrated data for 1951–2014. Metrics are averaged across spatial locations (pairs of locations for spatial correlation). Best (smallest) values for each combination of region, variable, and metric are highlighted.
Southwest Southeast
SPQR CCA QM SPQR CCA QM
TMAX Wasserstein distance 0.2673 0.1223 0.1057 0.0747 1.2413 1.2391
0.95 quantile 0.5617 0.3860 0.3416 0.4488 1.6876 1.6054
Lag-1 autocorrelation 0.0399 0.0764 0.0357 0.0336 0.0478 0.0297
Spatial correlation 0.0540 0.0834 0.0709 0.0188 0.0669 0.0579
PRCP Wasserstein distance 0.0747 0.4083 0.1537 0.2174 0.6851 0.5499
0.95 quantile 0.9172 0.7263 0.4264 2.2132 2.6586 2.5929
Proportion of zeros 0.0452 0.2624 0.2172 0.0340 0.1330 0.1722
Lag-1 autocorrelation 0.0921 0.1461 0.1138 0.0551 0.0878 0.0751
Spatial correlation 0.0851 0.1951 0.1016 0.0463 0.1334 0.1162
Cross correlation 0.0005 0.3258 0.3258 0.0004 0.0969 0.0969

### v3

@@ 4.3 Comparison of SPECD with CCA and QM
SPECD QM CCA SPECD QM CCA
TMAX Wass. distance 0.120 (0.035) 0.106 (0.020) 0.122 (0.051) 0.125 (0.029) 1.239 (1.162) 1.241 (1.163)
0.95 quantile 0.346 (0.478) 0.233 (0.328) 0.288 (0.326) 0.391 (0.511) 1.122 (1.596) 1.220 (1.648)
Spatial corr. 0.027 (0.034) 0.037 (0.044) 0.060 (0.047) 0.021 (0.026) 0.049 (0.071) 0.052 (0.072)
PRCP Wass. distance 0.075 (0.030) 0.154 (0.030) 0.408 (0.107) 0.216 (0.097) 0.550 (0.112) 0.685 (0.168)
0.95 quantile 0.741 (0.990) 0.318 (0.394) 0.546 (0.674) 1.921 (2.275) 1.983 (2.518) 2.096 (2.562)
Prop. of zeros 0.031 (0.026) 0.186 (0.119) 0.231 (0.128) 0.021 (0.023) 0.143 (0.107) 0.111 (0.085)
Spatial corr. 0.076 (0.094) 0.076 (0.099) 0.144 (0.088) 0.043 (0.053) 0.079 (0.101) 0.097 (0.112)
Cross corr. 0.001 (0.001) 0.307 (0.110) 0.307 (0.110) 0.000 (0.000) 0.082 (0.058) 0.082 (0.058)
(a) In-sample predictive accuracy (standard errors) for training period (1951–2000).
Southwest Southeast
SPECD QM CCA SPECD QM CCA
TMAX Wass. distance 0.328 (0.079) 0.307 (0.063) 0.927 (0.096) 0.898 (0.139) 1.674 (0.890) 1.345 (1.106)
0.95 quantile 0.912 (1.061) 0.946 (1.085) 0.749 (0.777) 1.022 (1.103) 1.767 (1.814) 1.276 (1.753)
Spatial corr. 0.037 (0.050) 0.037 (0.049) 0.042 (0.052) 0.036 (0.048) 0.057 (0.080) 0.058 (0.082)
PRCP Wass. distance 0.146 (0.060) 0.126 (0.026) 0.543 (0.154) 0.233 (0.050) 0.478 (0.143) 0.857 (0.235)
0.95 quantile 1.302 (1.604) 1.276 (1.650) 1.927 (1.530) 4.040 (5.057) 4.660 (5.791) 4.020 (3.442)
Prop. of zeros 0.103 (0.107) 0.171 (0.215) 0.176 (0.175) 0.086 (0.066) 0.166 (0.194) 0.084 (0.094)
Spatial corr. 0.128 (0.161) 0.117 (0.151) 0.096 (0.124) 0.080 (0.101) 0.099 (0.129) 0.091 (0.109)
Cross corr. 0.001 (0.001) 0.298 (0.121) 0.298 (0.121) 0.001 (0.001) 0.086 (0.079) 0.086 (0.079)
(b) Out-of-sample predictive accuracy (standard errors) for validation period (2001–2014).
Table 1: Predictive accuracy (with standard errors) in the density correction of TMAX and PRCP based on the SPECD model, compared with canonical correlation analysis (CCA) and quantile mapping (QM). Other than the Wasserstein distance metric, the remaining values correspond to the MAE between observed and calibrated data. Metrics are averaged across spatial locations (pairs of locations for spatial correlation). Best (smallest) values for each combination of region, variable, and metric are highlighted.
@@
To evaluate the performance of the SPECD density correction approach compared to CCA and QM, we provide metrics for predictive accuracy based on the three methods in Table 1. Marginal fits for TMAX and the transformed PRCP are summarized using the Wasserstein distance, while for the remainder of the metrics provided in this section, we evaluate the MAE. In each case, smaller values indicate more accurate predictions. Standard errors for the estimates are provided in parenthesis. For the training period (Table 1(a)), SPECD outperforms the competing methods for the SE across all metrics. In the more challenging SW region, QM is competitive against SPECD; it is often better at marginal metrics, whereas SPECD is better at spatial and cross correlations. For the validation period (Table 1(b)), all three methods are competitive in specific areas. In the SE, CCA is better at calibrating the lower and upper tails of PRCP, but SPECD is better at the remaining metrics. In the SW, there is no clear winner among the three methods; while SPECD has the highest MAE for spatial correlation of PRCP, it has lower error than one or both of its competitors across the remaining metrics.
@@
Overall, the SPECD calibrated distributions of TMAX and PRCP have the same characteristics as their observed data counterparts, for statistics of interest. In particular, it is able to very accurately capture the relationship between the two variables, which is crucial as they jointly affect several quantities of practical interest, including droughts, fire risk, and streamflow. Despite the differences in the spatial structure in the training and validation periods, it has competitive out-of-sample predictive accuracy, and excellent in-sample predictive accuracy.
@@ 5 Discussion
In this paper, we develop a new conditional density estimator for the simultaneous density correction of spatial fields of PRCP and TMAX. A key goal of our study is to maintain the cross correlation between the two variables during density correction. SPECD leverages the Vecchia approximation to maintain spatiotemporal and inter-variable dependencies, and density estimation is carried out using the deep learning based SPQR method. Our approach is trained across two spatial fields in the Southeast and Southwest US using historical data from 1951–2000, and tested using data from 2001–2014 for the same region. SPECD is able to outperform two competing methods (QM and CCA) in most scenarios, and the joint distribution characteristics of the observational data is preserved in the calibrated data.
@@
SPECD is highly scalable, since the Vecchia approximation allows for fitting models at different locations and months in parallel. Additionally, it ensures that the computational cost is linear in the number of grid cells. In our current implementation, however, we fit separate models for each month to capture finer seasonal effects. Future work will focus on consolidating the monthly models into a single annual model to reduce computational cost and improve estimation (since significantly more data will be available to train the annual model). We anticipate this can be achieved by including time-based covariates to account for trends and seasonality. Additionally, we plan to extend SPQR to support a bivariate response variable, which could then be used to directly model the joint distribution of TMAX and PRCP without needing to decompose it into a product of a conditional and a marginal distribution. Finally, we aim to explore the recently developed semi-parametric quantile regression for extremes framework (SPQRx; Majumder and Richards, 2025), as a replacement for SPQR. SPQRx is designed for modeling conditional densities with heavy tails, and will likely alleviate some of the issues with modeling the upper tails of TMAX and PRCP, and being more robust to distribution shifts than SPECD based on SPQR. The ultimate goal of these anticipated extensions is to provide future projections of TMAX and PRCP across the entire contiguous US.
Data availability
Preprocessed data and R Code for SPECD is available on GitHub at https://github.com/reetamm/SPECD. Additionally, the raw data as well as Python scripts for implementing the competing approaches (QM and CCE) are available on GitHub at https://github.com/fsq0511/NFbias.
Acknowledgments
@@
This work was supported by a grant from the National Science Foundation (DMS 2152887, CBET 2151651). Emily Hector was also supported by an Internationalization Seed Grant award from the Office of Global Engagement at North Carolina State University.
Appendix A Extension of SPECD to spatiotemporal data
Algorithm 3 Spatiotemporal density correction for $Y_{1}=$ TMAX and $Y_{2}=$ PRCP
Data $\mbox{\bf Y},X$
for $l=1,\ldots,n$ do
  $\mbox{\bf Y}_{(1lt)}=\{Y_{1l,t-1},Y_{1,l-1,t},\ldots,Y_{1,l-m,t}\}$
  procedure SPQR for TMAX and PRCP
   Estimate $F_{1l}$ and $Q_{1l}$ for $Y_{1lt}|\mbox{\bf Y}_{(1lt)},X$ $\triangleright$ Estimation step
   Estimate $F_{2l}$ and $Q_{2l}$ for $Y_{2lt}|\mbox{\bf Y}_{(2lt)},X$ $\triangleright$ Estimation step
  end procedure
  procedure TMAX density correction
   $u_{1lt}\leftarrow F_{1l}(Y_{1lt}|\mbox{\bf Y}_{(1lt)},X=0)$ $\triangleright$ Projection step
   for $t=2,\ldots,T$ do
     $Y_{1lt}^{*}\leftarrow Q_{1l}(u_{1lt}|\mbox{\bf Y}_{(1lt)}^{*},X=1)$ $\triangleright$ Calibration step
   end for
  end procedure
  procedure PRCP density correction
   $u_{2lt}\leftarrow F_{2l}(Y_{2lt}|\mbox{\bf Y}_{(2lt)},X=0)$ $\triangleright$ Projection step
   for t = 2,…,T do
     $Y_{2lt}^{*}\leftarrow Q_{2l}(u_{2lt}|\mbox{\bf Y}_{(2lt)}^{*},X=1)$ $\triangleright$ Calibration step
   end for
  end procedure
end for
We extend the model from Section 2.4 in the main text to incorporate temporal information. Let $Y_{1lt}$ and $Y_{2lt}$ denote the response for TMAX and PRCP at location $l\in\{1,\ldots,n\}$ and time $t\in\{1,...,T\}$. We order the $p=2nT$ variables as $(Y_{111},\ldots,Y_{11T},Y_{211},\ldots,Y_{21T},\ldots,Y_{1n1},\ldots,Y_{1nT},\\ Y_{2n1},\ldots,Y_{2nT})$. The SPECD model applies to multivariate time series data after key assumptions of Markovian and stationarity dependence structure. The Markovian structure states that the conditional quantile function for $Y_{jlt},j=1,2,$ given all observation before time $t$ depends only on data from the previous $T^{\prime}<T$ time steps; the stationarity assumption states that the conditional quantile functions are invariant to $t$.
We assume the conditional quantile function to be first order Markov in time, i.e., $T^{\prime}=1$. This assumption is reasonable given the selected are from the Southeast and Southwest US in regions where pronounced seasonality in precipitation is absent (Petersen et al., 2012). This implies that the Vecchia neighbor set for $Y_{1lt}$ is $\mbox{\bf Y}_{(1lt)}=\{Y_{1l,t-1},Y_{1,l-1,t},\ldots,Y_{11t}\}$, and for $Y_{2lt}$ is $\mbox{\bf Y}_{(2lt)}=\{Y_{2l,t-1},Y_{1l,t-1},Y_{1,l,t},\ldots,Y_{1,l-m,t},Y_{2,l-1,t},\ldots,Y_{2,l-m,t}\}$. These are straightforward extensions of the Vecchia neighbor sets presented in Section 2.4, with TMAX on day $t$ being additionally depending on TMAX on day t1t-1, and PRCP on day $t$ additionally depending on both PRCP and TMAX from day t1t-1. The conditional QF $Q$, CDF $F$, and calibration function $C$, for $\mbox{\bf Y}_{lt}$ are defined identically to Section 2.3; these do not depend on $t$ due to the stationarity assumption. Algorithm 3 details the density correction methodology for spatiotemporal data.
Appendix B QM and CCA implementation details
This section briefly outlines QM and CCA methodology that is employed to bias correct the GCM data. We refer readers to Bhowmik et al. (2017) and Seo et al. (2019) for further details on bias correction using QM and CCA.
B.1 Quantile mapping
Our QM implementation first sorts each time series (GCM and observation) individually in ascending order. It then applies a simple linear regression on the sorted data, with the observational data serving as the response, and the GCM data the covariate. Once the linear regression models have been fitted for each variable at each grid point, bias correction is carried out by substituting the raw (unsorted) GCM time series into the fitted regression equation. The steps are outlined in Algorithm 4.
Algorithm 4 Quantile mapping for TMAX and PRCP
Observational TMAX ($\mbox{\bf Y}_{1}$) and PRCP ($\mbox{\bf Y}_{2}$) data
GCM TMAX ($\mbox{\bf X}_{1}$) and PRCP ($\mbox{\bf X}_{2}$) data
for location $l=1,\ldots,n$ do
  procedure Asynchronous regression for TMAX and PRCP
   Obtain $Y_{1l}^{\prime},Y_{2l}^{\prime},X_{1l}^{\prime},X_{2l}^{\prime}$ by sorting $Y_{1l},Y_{2l},X_{1l},X_{2l}$ individually $\triangleright$ Sorting step
   Fit simple linear regression models $Y_{1l}^{\prime}\sim X_{1l}^{\prime}$ and $Y_{2l^{\prime}}\sim X_{2l}^{\prime}$ $\triangleright$ QM step
  end procedure
  procedure TMAX and PRCP bias correction
[…]
  end procedure
end for
B.2 Canonical correlation analysis
Canonical correlation is employed in bias correction to address one of the common issues that arise with asynchronous regression. Assume $\mbox{\bf Y}_{l}$ is a bivariate vector of observed TMAX and PRCP at location $l$. QM orders each component of this bivariate vector independent of the other; a similar operation is carried out for $\mbox{\bf X}_{l}$, the bivariate vector of GCM data. This affects the paired nature of the two meteorological variables, and there is no guarantee that the cross-correlation between TMAX and PRCP will be preserved in the bias-corrected output. Ideally, the bivariate vectors should be ordered in a manner that does not break the pairs, and bias correction carried out afterwards.
Our CCA approach begins by fitting separate bivariate Normal distributions to $\mbox{\bf Y}_{l}$ and $\mbox{\bf X}_{l}$. Each bivariate vector is then ordered based on their joint CDF, ensuring that (TMAX, PRCP) remains paired afterwards. The sorted variables are standardized using mean-variance transformations for each marginal distribution. CCA is carried out on the standardized data, $\mbox{\bf Y}_{l}^{\prime}$ and $\mbox{\bf X}_{l}^{\prime}$:
$U_{l}$ $=\mbox{\bf X}_{l}^{\prime}A_{l},$
$V_{l}$ $=\mbox{\bf Y}_{l}^{\prime}B_{l},$
Where $U_{l},V_{l}$ are rotated components of $\mbox{\bf Y}_{l},\mbox{\bf X}_{l}$ with maximum (canonical) correlation $R_{l}$, and $A_{l}$ and $B_{l}$ are the associated eigenvectors. The predicted $\hat{U}_{l}$ is obtained by plugging in $\mbox{\bf X}_{l}^{\prime}$, and predictions of $\hat{V}_{l}$ is obtained as follows:
$\hat{V}_{l}=R_{l}+(\hat{U}_{l}+\sqrt{1-R_{l}^{2}}).$
Next, we take the component-wise reciprocal of $B_{l}$ via an operation denoted as inv($B_{l}$), and use it to obtain a bias-corrected bivariate vector:
$\hat{\mbox{\bf Y}}_{l}=\hat{V}_{l}\cdot\mbox{inv}(B_{l}).$
Finally, the components of $\hat{\mbox{\bf Y}}_{l}$ are rescaled to their original units, and the sorting is reversed to obtain a time series of bias-corrected TMAX and PRCP.
Algorithm 5 Canonical correlation analysis for TMAX and PRCP
Observational TMAX ($\mbox{\bf Y}_{1}$) and PRCP ($\mbox{\bf Y}_{2}$) data, $\mbox{\bf Y}=(\mbox{\bf Y}_{1},\mbox{\bf Y}_{2})^{T}$
GCM TMAX ($\mbox{\bf X}_{1}$) and PRCP ($\mbox{\bf X}_{2}$) data, $\mbox{\bf X}=(\mbox{\bf X}_{1},\mbox{\bf X}_{2})^{T}$
for location $l=1,\ldots,n$ do
  procedure Asynchronous regression for TMAX and PRCP
   Fit bivariate Normal distribution to $\mbox{\bf X}_{l}$ and sort by joint CDF $\triangleright$ Sorting step
   Fit bivariate Normal distribution to $\mbox{\bf Y}_{l}$ and sort by joint CDF $\triangleright$ Sorting step
   Standardize $\mbox{\bf X}_{l}$ and $\mbox{\bf Y}_{l}$ marginally to obtain $\mbox{\bf X}_{l}^{\prime}$ and $\mbox{\bf Y}_{l}^{\prime}$
   Apply CCA to obtain rotated components $U_{l}$ and $V_{l}$, eigenvectors $A_{l}$ and $B_{l}$, and $R_{l}$, the canonical correlation between $U_{l}$ and $V_{l}$ $\triangleright$ CCA step
  end procedure
  procedure TMAX and PRCP bias correction
   Obtain sorted predictions for the rotated components:
   $\hat{U}_{l}\leftarrow\mbox{\bf X}_{l}^{\prime}A_{l}$
   $\hat{V}_{l}\leftarrow R_{l}+\hat{U}_{l}\sqrt{1+R_{l}^{2}}$
   Obtain sorted and standardized predictions $\hat{\mbox{\bf Y}}_{l}^{\prime}\leftarrow\hat{V}_{l}\cdot\mbox{inv}(B_{l})$ $\triangleright$ Bias correction step
   Scale and unsort to obtain bias-corrected daily records $\hat{\mbox{\bf Y}}_{l}$
  end procedure
end for
Appendix C Additional numerical study results
This section contains additional results for the numerical study presented in Section 3 of the main text. Table 2 describes the numerical results corresponding to Figure 4. Each value in the tables corresponds to a metric averaged over 100 replicates of the simulation study. Values in the parentheses represent the standard errors of the estimates.
QM CCE SPECD1 SPECD2
TMAX Wass. distance 1.7888 (0.1267) 2.0463 (0.5496) 0.0993 (0.0112) 0.2697 (0.0216)
0.95 quantile 1.811 (0.1339) 2.3730 (0.7139) 0.1320 (0.0328) 0.5840 (0.0589)
Spatial corr. 0.4831 (0.0098) 0.7987 (0.0072) 0.4723 (0.0112) 0.1411 (0.0208)
PRCP Wass. distance 0.1037 (0.0034) 0.1123 (0.0014) 0.0265 (0.0018) 0.0680 (0.0080)
0.95 quantile 0.1819 (0.0167) 0.1028 (0.0167) 0.0725 (0.0132) 0.2462 (0.0693)
Spatial corr. 0.4770 (0.0123) 0.8357 (0.0107) 0.3787 (0.0230) 0.2607 (0.0199)
Prop. of zeros 0.2050 (0.0040) 0.2069 (0.0020) 0.0261 (0.0036) 0.0363 (0.0077)
Cross corr. 0.3364 (0.0195) 0.3031 (0.0052) 0.0005 (2.89e-05) 0.0004 (2.39e-04)
(a) In-sample predictive accuracy (standard errors) for training period (1951–2000).
QM CCE SPECD1 SPECD2
TMAX Wass. distance 1.7917 (0.1414) 2.0812 (0.5655) 0.2681 (0.0436) 0.3552 (0.0378)
0.95 quantile 1.8215 (0.1632) 2.3480 (0.7162) 0.3398 (0.0749) 0.6236 (0.1232)
Spatial corr. 0.4839 (0.0193) 0.7955 (0.0164) 0.4728 (0.0212) 0.1517 (0.0291)
PRCP Wass. distance 0.1067 (0.0089) 0.0785 (0.0033) 0.0471 (0.0089) 0.0801 (0.0175)
0.95 quantile 0.1894 (0.0395) 0.1381 (0.0332) 0.1375 (0.0321) 0.2949 (0.1060)
Spatial corr. 0.4750 (0.0273) 0.8326 (0.0246) 0.3970 (0.0324) 0.2514 (0.0459)
Prop. of zeros 0.2041 (0.0151) 0.1683 (0.0070) 0.0349 (0.0108) 0.0393 (0.0101)
Cross corr. 0.3376 (0.0326) 0.4032 (0.0113) 0.0007 (0.0003) 0.0012 (0.0007)
(b) Out-of-sample predictive accuracy (standard errors) for validation period (2001–2014).
@@ Appendix C Additional numerical study results
Table 2: Predictive accuracy estimates (standard errors) in the density correction of TMAX and PRCP based on QM, CCA, and SPECD, based on a simulation study with 100 replicates. Other than the Wasserstein distance metric, the remaining values correspond to the MAE between observed and calibrated data. Metrics are averaged across spatial locations (pairs of locations for spatial correlation). Best (smallest) values for each combination of region, variable, and metric are highlighted.
Appendix D Non-stationarity in the spatial dependence
Figure 11: Comparison of pairwise spatial correlations for TMAX in the SW region between the training (1951–2000) and validation (2001–2014) periods.
Figure 12: Comparison of pairwise spatial correlations for PRCP in the SW region between the training (1951–2000) and validation (2001–2014) periods.
Figure 13: Comparison of pairwise spatial correlations for TMAX in the SE region between the training (1951–2000) and validation (2001–2014) periods.
Figure 14: Comparison of pairwise spatial correlations for PRCP in the SE region between the training (1951–2000) and validation (2001–2014) periods.
In this section, we aim to study a shift in the spatial correlations between the training and validation periods of our study. We achieve this by comparing the pairwise spatial correlations for each combination of month, variable, region, and data source (model vs. observed). Ideally, for any of these combinations, we would like the spatial distribution to remain unchanged over time if we want to accurately assess the out-of-sample performance of the density correction methods. Figures 1114 plot the pairwise spatial correlations with the training and validation periods on the two axes. Model data is represented in red, and observational data is represented in black. Deviation of points from the diagonal could indicate one of the following (non-exhaustive) scenarios:
  1. 1.
    Low variability, bias from one source: Both data sources show low variability in the their estimates for the two periods, but one of the data sources shows bias; e.g., the spatial correlations in the training period is higher/lower than the spatial correlations in the validation period.
  2. 2.
    Low variability, bias in both sources: Both data sources show low variability and both of them show bias. The bias may be in the same or opposite directions; e.g., the correlations in the model data increase while those in the observational data decrease.
  3. 3.
    High variability, low bias: Both data sources show high variability, i.e., the spatial correlation estimates differ noticeably in the two time periods. However, bias is low, i.e., the differences in the spatial correlations from the two time periods are centered at zero.
  4. 4.
    High variability, bias from at least one source: Both data sources show high variability in their estimates. Additionally, at least one of the variables has bias. If both variables have bias, they may be in the same direction or in opposite directions.
Figures 11 and 13 plot the spatial correlations in TMAX for the SW and SE regions. These plots have low variability, and largely low bias. Some of the months correspond to Scenario 1 (e.g., August and December plots for the SE) or Scenario 2 (e.g., April and July in the SW).
Figures 12 and 14 plot the spatial correlations in PRCP for the SW and SE regions. These plots have higher variability, likely due to the presence of zeros in the data. Additionally, the SW has higher variability than the SE. Plots for the SE mostly correspond to Scenario 3, but some plots for the SW do demonstrate behavior consistent with Scenario 4 (e.g., February and October).
Overall, we also notice more deviations from the diagonal that correspond to relatively low correlation values, and stronger agreement when the correlations are the highest. These results affect the calibration based on SPECD across all its metrics, not just the spatial correlations. This is because spatial information is encoded within the model by the Vecchia approximation, whereby values at nearby grid cells are used as covariates. In particular, we would expect to see better calibration for TMAX than for PRCP, and for the SE compared to the SW.
Appendix E Additional seasonal results
This section provides additional seasonal and monthly calibration metrics of PRCP and TMAX. Figure 15 plots seasonal densities of the two variables for both regions, for the training and validation periods. The seasons are defined as Spring (Jan–Mar), Summer (Apr–Jun), Fall (Jul–Sep), and Winter (Oct–Dec). The density of TMAX is well calibrated for both the training and validation periods. The PRCP densities are also well calibrated, though there are a few deviations around the extremes for the validation period. In particular, the proportion of zeros is overestimated for both regions in several of the months.
Figure 16 plots the monthly 0.95 quantiles of both variables, and the proportion of zeros for PRCP; these are analogous to the monthly mean plots (Figure 10) in the main text. We note that the TMAX quantiles are well calibrated for both regions and both time periods (training and validation), with a little more error in the validation period compared to the training period. Similar to the corresponding plots for the means, the distribution of the quantiles are quite similar during the two time periods, and therefore are calibrated quite well. The proportion of zeros for the training period are slightly underestimated in both regions, since there is a massive difference between the proportions in the model and observational data; despite this, they are overall well calibrated. For the validation period, several of the months have lower proportion of zeros in the observational data compared to the historical period. In such situations, the calibrated data can overestimate the proportion. However, the calibration is in general adequate. Finally, for the upper quantiles of PRCP, the calibration for the SW is adequate for both the training and validation periods. However, for the SE, SPECD struggles to calibrate the quantiles for both the training and validation periods. This is similar behavior to their monthly means, and demonstrates an area where SPECD can be improved upon.
(a) Seasonal density plots for training period of 1951–2000.
(b) Seasonal density plots for validation period of 2001–2014.
Figure 15: Seasonal density plots for TMAX and PRCP data in the Southwest (SW) and Southeast (SE).
(a) Upper quantiles and proportion of zeros for training period of 1951–2000.
(b) Upper quantiles and proportion of zeros for validation period of 2001–2014.
Figure 16: Monthly 0.95 quantiles (TMAX and PRCP) and proportion of zeros (PRCP) for the Southwest (SW) and Southeast (SE), pooled across all 25 locations.